Find the market equilibrium if P=-40+Q and P=80-Qd
1
Expert's answer
2020-04-28T07:47:29-0400

Equilibrium occurs at the point where the quantity demanded is equal to the quantity supplied.


Equating the demand curve to the supply curve and solving for Q:



"-40 + Q = 80- Q"

"2Q = 120"

"Q^* = \dfrac{120}{2}""\boxed{\color{blue}{Q^* = 60}}"


The equilibrium price is:



"P^* = -40 + 60"

"\boxed{\color{blue}{P^* = \$20}}"